Sodium-24 has a half-life of 15 hours. How much sodium-24 will remain in an 18.0g sample after 69 hours


1
Expert's answer
2021-03-29T09:02:17-0400
"m=m_0\cdot 2^{-\frac{t}{T}}\\m=18\cdot 2^{-\frac{69}{15}}=0.742\ g"